### v2

# VidHal: Benchmarking Temporal Hallucinations in Vision LLMs

Abstract
Vision Large Language Models (VLLMs) are widely acknowledged to be prone to hallucinations. Existing research addressing this problem has primarily been confined to image inputs, with limited exploration of video-based hallucinations. Furthermore, current evaluation methods fail to capture nuanced errors in generated responses, which are often exacerbated by the rich spatiotemporal dynamics of videos. To address this, we introduce VidHal, a benchmark specially designed to evaluate video-based hallucinations in VLLMs. VidHal is constructed by bootstrapping video instances across a wide range of common temporal aspects. A defining feature of our benchmark lies in the careful creation of captions which represent varying levels of hallucination associated with each video. To enable fine-grained evaluation, we propose a novel caption ordering task requiring VLLMs to rank captions by hallucinatory extent. We conduct extensive experiments on VidHal and comprehensively evaluate a broad selection of models. Our results uncover significant limitations in existing VLLMs regarding hallucination generation. Through our benchmark, we aim to inspire further research on I) holistic understanding of VLLM capabilities, particularly regarding hallucination, and II) extensive development of advanced VLLMs to alleviate this problem. Our VidHal dataset and evaluation code are publicly available at https://github.com/Lookuz/VidHal.
1 Introduction
Building on the advancements of Large Language Models (LLMs), Vision LLMs (VLLMs) have recently gained significant attention. Models such as LLaVA [33, 34] have shown impressive performance across various visual understanding tasks involving both images and videos. Despite their potential, VLLMs are notably prone to hallucinations, where generated responses that appear to be plausible contradict the visual context [1, 57]. This problem significantly compromises the reliability of VLLMs, hindering their practical use in real-world applications.
Figure 1: Multiple-Choice Question Answering (MCQA) performance of representative VLLMs on our VidHal benchmark. (Left) Overall ranking of VLLMs. (Right) Detailed accuracy results pertaining to each temporal aspect, wherein higher scores indicate fewer hallucinations.
To tackle this challenge, some methods propose to leverage post-hoc techniques such as contrastive decoding [23, 75, 12, 76] and attention calibration [17, 40, 38, 64, 15, 69, 56]. Other efforts have been devoted to the evaluation of hallucinations in VLLMs. For example, CHAIR [46] initially studies object-based hallucination evaluation with the aid of the image captioning task. Subsequent studies [29, 37, 21, 10] instead harness paired ⟨positive, hallucinatory⟩ questions to probe such hallucinations. Additionally, MMHalBench [49] and AMBER [51] expand beyond object-based evaluations by constructing benchmarks that cover attribute and relationship hallucinations.
Unlike their image-based counterparts, video hallucinations pose unique challenges primarily due to the intricate spatiotemporal dynamics of videos [28, 44, 5, 13, 39, 41]. In particular, video-specific temporal aspects, such as movement direction and chronological order of events, are especially concerning for video-based VLLMs. Furthermore, the richness of video content necessitates a finer-grained understanding, making VLLMs more vulnerable to nuanced hallucinations. Nonetheless, to the best of our knowledge, video-based hallucinations remain underexplored in the existing literature.
To address this research gap, we present VidHal, a benchmark specifically designed to evaluate video-based hallucinations of VLLMs. VidHal features videos that comprehensively cover a broad range of temporal aspects, such as entity actions and sequence of events. Each video is automatically annotated with multiple captions exhibiting varying levels of aspect-specific hallucinations, capturing both subtle and significant discrepancies. In addition, we perform detailed human validation to ensure the robustness and reliability of our annotation process. An additional motivation stems from the limited metrics for quantifying hallucinations in VLLMs. To capture fine-grained hallucinatory errors of these models, we propose a unique caption ordering task that requires models to rank captions by hallucination levels. This consequently leads to a ranking-based NDCG metric and an MCQA accuracy metric, both are distinct from prior ones and specifically tailored to evaluate nuanced hallucinations in video-based VLLMs.
Using our VidHal dataset, we benchmark thirteen VLLMs including both open-sourced and proprietary models, with abstracted results summarized in Figure 1. Through these extensive experiments, we identify limitations in nuanced video understanding among all evaluated VLLMs. Specifically, our findings reveal that existing VLLMs struggle to differentiate between captions with varying levels of hallucination. This deficiency is particularly evident when evaluating video-specific aspects, such as Direction and Order, as illustrated in Figure 1, indicating substantial room for improvement in current video-based VLLMs. Besides, we also notice that proprietary models, e.g., GPT-4o [42], frequently outperform open-source models by significant margins.
Overall, the contributions of this work are three-fold:
  • We present VidHal, a benchmark dataset dedicated to video-based hallucination evaluation of VLLMs. Our dataset is distinguished by I) video instances sourced from public video understanding datasets encompassing a diverse range of temporal concepts and II) captions with varying levels of hallucination.
  • We introduce a novel evaluation task of caption ordering along with two metrics designed to evaluate fine-grained hallucination generation in existing VLLMs.
  • We conduct extensive experiments on VidHal with a variety of VLLMs, uncovering limitations in their fine-grained video reasoning abilities, particularly in their tendency to generate hallucinations.
2 Related Work
2.1 Vision Large Language Models
The emergence of powerful LLMs has advanced the development of VLLMs [33, 34, 26, 9, 60, 61, 59]. Typical methods in this category include LLaVA [33], mPLUG-Owl [61, 59, 60], InstructBLIP [9], and MiniGPT-4 [73]. These VLLMs rely on aligning vision encoders with LLMs using connective modules such as Q-Former [9, 25, 26, 65, 8] or MLPs [33, 34, 48] with the instruction tuning stage. Recent methods have extended visual inputs from images to (long) videos, delivering impressive joint spatial-temporal reasoning capabilities. For instance, VideoLLaMA2 [8] enhances the LLaMA model with video understanding capabilities through a Spatial-Temporal Convolution (STC) module. LLaVA-NeXT-Video [35, 66] presents an AnyRes approach that enables reasoning with long videos.
2.2 Hallucinations in VLLMs
Despite their impressive performance on visual reasoning benchmarks, current VLLMs remain notoriously susceptible to hallucinations [19, 38, 74, 7]. A common demonstration is that the generated responses contain information which is inconsistent with the visual content [1, 36, 63, 55]. Most approaches address the hallucination problem with post-hoc techniques. For example, LURE [71] and Woodpecker [62] develop pipelines that assist VLLMs in revising their responses using expert models. To reduce bias from unimodal and statistical priors, contrastive decoding methods, such as VCD [23] and M3ID [12], along with attention calibration techniques like OPERA [17] are employed to refine token predictions. Building on the success of reinforcement learning for preference optimization in LLM development [43], HA-DPO [67], POVID [70] and CSR [72] adopt this paradigm to fine-tune VLLMs, yielding outputs with fewer hallucinations.
Figure 2: Overview of our VidHal benchmark construction pipeline. Using direction as an example from the five selected aspects, we begin by sourcing relevant video instances from existing datasets. Next, the anchor (positive) caption is generated from the original video metadata. Finally, GPT-4o is employed to generate hallucinatory captions at varying levels.
2.3 Video Reasoning Benchmarks
The rise of video-based VLLMs has driven the development of numerous video benchmarks. Notable examples, such as SEEDBench [24], VideoBench [41], MVBench [28], and VideoMME [13], focus on dynamic events requiring temporal reasoning beyond individual frames. However, these benchmarks often lack diversity in reasoning tasks and visual concepts. To address this, AutoEval-Video [5] and Perception Test [44] introduce complex reasoning tasks such as counterfactual and explanatory reasoning, while TempCompass [39] expands temporal concept coverage. Several benchmarks [29, 51, 49, 21, 31, 20, 53, 68, 7, 4, 54, 50, 6] have been constructed to quantify visual hallucinations, primarily targeting object-based hallucinations in images. HallusionBench [16], VideoCon [2], and Vript [58] provides partial coverage of video-based hallucinations, while VideoHallucer [52] introduces a benchmark for hallucination detection in videos. However, these benchmarks provide limited coverage of spatio-temporal concepts, focusing on conventional aspects like actions and objects while neglecting other video-centric elements such as direction. Additionally, their evaluation strategies primarily follow image-based approaches, which we argue are less effective in capturing nuanced, video-specific hallucinations.
3 VidHal Dataset Construction
We introduce VidHal, a unique video-language benchmark designed to evaluate hallucinations of Video-LLMs in a comprehensive manner. As depicted in Figure 2, VidHal comprises of video instances which span a diverse spectrum of temporal aspects, including previously unexplored aspects such as directional movement. In contrast to previous studies on video hallucination evaluation [58, 52, 2], VidHal incorporates multiple hallucinated captions per video, enabling the assessment of video hallucinations at multiple levels of granularity.
3.1 Temporal Hallucinations in Videos
Hallucinations in VLLMs occur when the model fabricates details in its responses that contradict the provided visual content. Compared to images, video hallucinations extend beyond static visual elements to include misperceptions of dynamic changes within scenes. We categorize these temporal hallucinations into two semantic levels:
Lexical Semantics (L-Sem) captures instances where VLLMs misinterpret words related to temporal features, including nouns referring to objects or attributes (e.g., misidentifying a color change from green to red as green to orange) and verbs describing actions (e.g., interpreting ”kicking a ball” as ”throwing a ball”).
Clause Semantics (C-Sem) encompasses errors involving event descriptions and their sequences, where the VLLM incorrectly predicts the order of events occurring in the video. For example, given sequentially occurring events $A$ and $B$ in a video, the model may perceive $B$ preceding $A$.
By addressing these two dimensions of video-based hallucinations, VidHal offers holistic coverage over the level of detail in which VLLMs may hallucinate.
3.2 Temporal Concept Selection
Prior research on hallucination evaluation for both images [29, 51, 46] and videos [52, 58, 16] has predominantly focused on common visual aspects such as action- and object-based hallucinations. However, video-based hallucinations may involve additional dynamic factors associated with spatio-temporal patterns, which are overlooked in these studies. In light of this, we propose to focus on the following five aspects to ensure comprehensive coverage of temporal concepts. Specifically, the first four aspects address hallucinations based on lexical semantics, while the fifth targets those based on clause-level semantics.
  • Attribute (L-Sem) describes the fine-grained characteristics and properties of objects or subjects in the video. We additionally categorize this aspect into sub-aspects of Size, Shape, Color, Count and State Change.
  • Object (L-Sem) relates to the interactions between objects and entities within the video. We further delineate this aspect into two fine-grained sub-aspects: Object Recognition, identifying the objects engaged in interactions, and Interaction Classification which concentrate on how these objects interact with other objects or subjects.
  • Action (L-Sem) refers to the movements and behaviours exhibited by entities in dynamic video scenes.
  • Direction (L-Sem) indicates the orientation and movement trajectory of subjects or objects in the video.
  • Event Order (C-Sem) represents the correct sequence of events in the video. During our collection, we retain videos that contain at least three distinct events.
We present an example that illustrates the direction aspect in Figure 2, with additional examples available in the supplementary material.
3.3 Hallucinatory Caption Generation
Based on the aspects in Section 3.2, we build our benchmark upon four public video understanding datasets: TempCompass [39], Perception Test [44], MVBench [28] and AutoEval-Video [5]. TempCompass and MVBench extensively cover all five temporal aspects, while Perception Test and AutoEval-Video highlights human-object interactions and attribute changes, respectively.
Existing hallucination benchmarks [29, 51] rely mostly on binary questions for evaluation, limiting their efficacy in detecting subtle video hallucinations, such as minor event inconsistencies. To address this issue, we advocate a novel evaluation protocol incorporating several carefully annotated captions. Specifically, each video will be annotated with a set of $M$ captions that reflect varying degrees of hallucination in VLLMs. Given the cost and labor intensity of manual annotation, we opt for automatic caption generation using a carefully designed pipeline illustrated in Figure 2.
Anchor Caption Generation. The video instances in VidHal are sourced from various public datasets, resulting in distinct associated metadata such as long-form captions in AutoEval-Video and question-answer pairs in MVBench. To ensure structure consistency and information granularity in the respective dataset description across all instances, we automatically generate an anchor caption for each video. Specifically, we input the metadata for each video $V^{i}$ into GPT-4o and prompt it to generate a concise and accurate description $y^{i}_{+}$ using the provided metadata information.
Hallucinatory Caption Generation. After obtaining the positive caption for each video instance, we augment the dataset with $M-1$ additional captions containing hallucinated content. For a given video instance $V^{i}$, we construct a set $\mathcal{Y}^{i}_{-}=\{y^{i,1}_{-},\cdots,y^{i,M-1}_{-}\}$ containing captions with different levels of hallucination based on the temporal concepts associated with it. Specifically, $y^{i,k}_{-}$ exhibits heavier hallucination than $y^{i,j}_{-}$ for $j<k$. We leverage GPT-4o to generate $\mathcal{Y}^{i}_{-}$ by combining the anchor caption $y^{i}_{+}$ and prompting it to create $y^{i,1}_{-},\cdots,y^{i,M-1}_{-}$ progressively in increasing levels of hallucination. The set of captions associated with $V^{i}$ is then defined as $\mathcal{Y}^{i}\leftarrow\{y^{i}_{+}\}\bigcup\mathcal{Y}^{i}_{-}$ consisting of both the anchor and hallucinatory captions.
Figure 3: Human agreement on hallucination levels in the VidHal dataset. (Left) Distribution of agreement ratios per video sample. (Right) Average agreement ratio for each temporal aspect, with an overall average of 87%.
3.4 Dataset Statistics and Human Validation
Our VidHal benchmark consists of a total of 1,000 video instances. Using our automatic annotation pipeline, each video instance is tagged with $M\mathbin{=}3$ captions. As shown in Table 1, our VidHal dataset stands out from other video understanding [24, 41, 28, 13] and hallucination benchmarks [2, 16, 32] in terms of two dimensions: I) VidHalencompasses a diverse range of video-centric temporal aspects; and II) We introduce a novel caption ordering task along with two tailored metrics to capture subtle hallucinations that are previously ignored by paired questions .
Dataset Temporal Aspects Task Formats Evaluation Metrics
Action Attribute Direction Object Order
Size Shape Color Count State-Change Recognition Interaction
[…]
Reasoning
SEEDBench [24] MCQA Accuracy
VideoBench [41] MCQA Accuracy
MVBench [28] MCQA Accuracy
Video-MME [13] MCQA Accuracy
[…]
Evaluation
Vript [58] Video Captioning F1 Score
Event Ordering Accuracy
VideoCon [2] VL Entailment ROC-AUC
HallusionBench [16] Y/N QA Accuracy
VidHal(Ours) MCQA Accuracy
Caption Ordering NDCG
Table 1: Comparison of our benchmark dataset with existing video-based reasoning and hallucination evaluation datasets. For datasets with multiple evaluation tasks, only those relevant to hallucination evaluation are included. VL Entailment denotes the task of video-language entailment, while Event Ordering prompts the model to determine the chronological sequence of scenes in a video.
To ensure the reliability of our generated captions at varying levels, we randomly selected 100 examples for human validation, where each sample is labeled by 15 annotators on average. Our human validation process focuses on verifying that the order of hallucinatory captions generated by our pipeline aligns with human judgment. Figure 3 reflects an overall agreement rate of 87% between our automatically generated hallucinatory captions and human annotators, indicating consistency between these two across all video-centric temporal aspects.
4 VidHal Evaluation Protocol
Figure 4: Visual illustration of relative caption ordering task in VidHal. The final ordering is parsed based on VLLM responses for each pair order queried.
Aiming to address the limitations of binary question-based benchmarks, we propose two evaluation tasks: multiple-choice question answering and a novel caption ordering task, detailed in Section 4.1. We also develop corresponding metrics to comprehensively measure hallucinations in video-based VLLMs, elaborated further in Section 4.2.
4.1 Evaluation Tasks
Multiple-Choice Question Answering (MCQA) assesses the model’s spatiotemporal understanding in a coarse-grained manner. Specifically, the model is provided with a video $V^{i}$ and its corresponding set of captions $\mathcal{Y}^{i}$ as answer options. The VLLM is then instructed to select the most appropriate caption for the video.
Caption Ordering evaluates a model’s visual reasoning from a nuanced granularity, instructing VLLMs to order the provided captions based on their hallucination level. Specifically, we design two caption ordering sub-tasks. The first, naive caption ordering, requires VLLMs to rank all captions at once. However, this sub-task can confuse several VLLMs due to its inherently challenging nature and the inferior instruction-following capabilities of some models. As a complement, we propose an additional sub-task, relative caption ordering, which decomposes the prior task into multiple paired caption ordering tasks. Since each paired ordering task is answered in isolation, the VLLM may produce a non-transitive, cyclic ranking. To circumvent this, we first query the model with consecutive caption pairs and prompt it with the final pair only if ambiguities arise in the initial sequence, as illustrated in Figure 4.
Notably, our relative caption ordering task is more challenging than previous binary questions. This complexity arises from certain paired questions in VidHal where both options are hallucinatory, making them harder to distinguish as opposed to ⟨positive, hallucinatory⟩ pairs.
Model Release Date Vision Encoder LLM #Params #Frames Accuracy NDCG
Naive Relative
Baseline
Random - - - - - 0.326 0.505 0.480
Open-Sourced Models
VideoChat [27] May 2023 EVA-CLIP-G Vicuna 7B 8 0.381 0.475 0.488
LLaMA-VID [30] Nov. 2023 EVA-CLIP-G Vicuna 7B 1fps 0.358 0.486 0.521
VideoChat2 (Vicuna) [28] Nov. 2023 UMT-L Vicuna 7B 16 0.426 0.486 0.577
VideoChat2 (Mistral) May 2024 UMT-L Mistral 7B 16 0.443 0.503 0.475
VideoChat2 (Phi) Jun. 2024 UMT-L Phi3 3.8B 16 0.514 0.626 0.612
mPLUG-Owl3 [59] Aug. 2024 SigLIP/SO400M Qwen2 7B 16 0.596 0.641 0.707
LLaVA-NeXT-Video (7B) [66] Apr. 2024 SigLIP/SO400M Vicuna 7B 32 0.509 0.518 0.620
LLaVA-NeXT-Video (32B) Apr. 2024 SigLIP/SO400M Qwen1.5 32B 32 0.663 0.641 0.747
VideoLLaMA2 (7B) [8] Jun. 2024 CLIP ViT-L/14 Mistral 7B 8 0.541 0.564 0.622
VideoLLaMA2 (72B) Jun. 2024 CLIP ViT-L/14 Qwen2 72B 8 0.647 0.787 0.760
Proprietary Models
GPT-4o [42] May 2024 - - - 1fps 0.772 0.840 0.826
Gemini-1.5 (Flash) [45] May 2024 - - - 1fps 0.657 0.738 0.745
Gemini-1.5 (Pro) May 2024 - - - 1fps 0.671 0.765 0.753
Table 2: Benchmark performance of VLLMs on our VidHal dataset. #Params refers to the number of parameters of the base LLM used. The best performance for each task is highlighted in bold for open-sourced models, and underlined for closed-sourced models.
4.2 Evaluation Metrics
Notations
For a particular video instance $V^{i}$, we define the ground truth caption order for $V^{i}$ to be $\mathcal{Y}^{i}_{*}=(y^{i}_{+},y^{i,1}_{-},\cdots,y^{i,M-1}_{-})$. Further let the $j^{th}$ element in this ordering be indexed as $\mathcal{Y}^{i,j}_{*}$.
MCQA
We employ the standard accuracy metric:
$\text{Accuracy}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[R_{MCQA}(V^{i},\mathcal{Y}^{i})=y^{i}_{+}\right],$ (1)
where $N$ is the number of video instances, $\mathbb{I}$ denotes the indicator function, and $R_{MCQA}(V^{i},\mathcal{Y}^{i})$ represents the best matched caption from $\mathcal{Y}^{i}$ for $V^{i}$ as predicted by a VLLM.
Caption Ranking
Inspired by metrics from the information retrieval domain [14], we adapt the well-established Normalized Discounted Cumulative Gain (NDCG) [18] for hallucination assessment in VidHal. Unlike previous metrics like POPE [29], our metric awards partial credit for correctly ordered caption pairs even when the optimal ranking is not achieved. As such, we expect the metric to effectively capture and distinguish both subtle and severe hallucinations generated by video-based VLLMs. Formally, we define our adapted NDCG metric as follows:
$\text{NDCG}=\frac{1}{N}\sum^{N}_{i=1}\frac{\text{DCG}_{i}-\text{rDCG}_{i}}{\text{iDCG}_{i}-\text{rDCG}_{i}},$ (2)
where $\text{DCG}_{i}$ is formulated as:
$\text{DCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)},$ (3)
and $\hat{y}^{i,j}$ represents $j^{th}$ caption in the ranked order predicted by the VLLM. The perfect ordering is achieved when $\hat{y}^{i,1}=y^{i}_{+}$ and $\{\hat{y}^{i,j}=y^{i,j-1}_{-}\}_{j=2\to M}$. To evaluate predicted caption orders relative to this ideal sequence, a relevance function $r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)$ is designed to assign higher scores to $\hat{y}^{i,j}$ with lower hallucinatory extent.
$r(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})=M+1-\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}),$ (4)
where $\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})$ denotes the position of $\hat{y}^{i,j}$ in $\mathcal{Y}^{i}_{*}$. Finally, $\text{DCG}_{i}$ is normalized to a range of $[0,1]$ using $\text{iDCG}_{i}$ and $\text{rDCG}_{i}$, with a score of 1 indicating perfect alignment of the predicted order with $\mathcal{Y}^{i}_{*}$. Specifically, these terms represent the maximum and minimum $\text{DCG}_{i}$ scores obtained from the optimal ordering $\mathcal{Y}^{i}_{*}$ and its reverse, respectively,
$\text{iDCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\mathcal{Y}^{i,j}_{*},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)},$ (5)
$\text{rDCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\mathcal{Y}^{i,M-j}_{*},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)}.$ (6)
5 Experiments
5.1 Experimental Settings
Models. We evaluated thirteen VLLMs from eight different model families, including six open-source models: VideoChat [27], LLaMA-VID [30], VideoLLaMA2 [8], VideoChat2 [28], mPLUG-Owl3 [59] and LLaVA-NeXT-Video [66], and two proprietary models: GPT-4o [42] and Gemini-1.5 [45]. These models represent a wide variety of architectural designs and training paradigms. Additionally, we included a random baseline that selects and ranks candidate options randomly.
Implementation Details. All experiments were conducted using four NVIDIA A100 40GB GPUs. The input captions in $\mathcal{Y}^{i}$ were presented in a randomized order using a fixed, predefined randomization seed across experiments. We adhered to the inference and model hyperparameters outlined in the respective original models, and employed greedy decoding during generation for a fair comparison.
5.2 Overall Results
Benchmark Results
We present the overall results of representative VLLMs in Table 2 across both MCQA and caption ordering tasks. We make three key observations from this table: I) Proprietary models demonstrate superior results compared to open-sourced models. In particular, GPT-4o achieves the best performance on all tasks, surpassing other models by significant margins. II) Larger VLLMs generally outperform smaller ones in both tasks. This result is supported by the comparison of different LLM bases for the VideoLLaMA2 and LLaVA-NeXT-Video models. III) The caption ordering task poses greater difficulty for current VLLMs than MCQA, evidenced by the larger performance margins between the VLLM models and the random baseline. Notably, VideoChat and VideoChat2 (Mistral) show slight to no improvement over the random baseline across both caption ordering tasks. This indicates that current VLLMs greatly suffer from poor fine-grained video understanding and are inclined to generate hallucinations.
Figure 5: Aspect-aware results of VLLMs for the (Left) naive and (Right) relative caption ordering task. The dotted lines represent the average NDCG scores across all models.
Aspect-aware Results
Figure 5 highlights the fine-grained, aspect-specific results. It can be observed that these VLLMs exhibit notably stronger performance on Action and Object aspects than the others. This is likely attributed to current visual instruction tuning data emphasizing object-centric recognition and coarse-grained activity classification, allowing models to leverage image-based priors for predictions. In contrast, these models struggle with fine-grained aspects related to direction and event order, which are especially unique to videos.
We further analyzed the distribution of results for the relative caption ranking task across sub-aspects of the Attribute and Object aspects in Figure 6. While VLLMs generally maintain consistent performance across Attribute sub-aspects, they demonstrated slightly less effectiveness in inference involving Count and Color. For the Object aspect, several models performed significantly worse in Interaction Classification than in Object Recognition. This highlights the need for greater emphasis on modeling object-based interactions to bridge the gap between object recognition and interaction understanding in VLLMs.
Figure 6: NDCG scores for the Attribute (Left) and Object (Right) sub-aspects in the relative caption ranking task.
5.3 Ablation Studies
VideoChat2 (Vicuna) mPLUG-Owl3 LLaVA-NeXT -Video (7B) VideoLLaMA2 (7B)
$P_{\text{TI}}\ (\%)$ 7.3 3.0 3.7 4.0
Table 3: Percentage of VidHal instances where final predicted orders contradict individual paired-caption orders.
Transitive Robustness
Despite improving VLLM stability in caption ordering, the relative caption ranking task may reveal transitive inconsistencies between the final caption order and individual paired orders. For instance, the predicted order $\mathcal{\hat{Y}}^{i}=[\hat{y}^{i,1},\hat{y}^{i,2},\hat{y}^{i,3}]$ may be inferred when the VLLM predicts $\hat{y}^{i,1}\prec\hat{y}^{i,2}$ and $\hat{y}^{i,2}\prec\hat{y}^{i,3}$, but it may predict $\hat{y}^{i,3}\prec\hat{y}^{i,1}$ when prompted to order these two captions, contradicting the initial order. We quantify such errors by measuring the proportion of instances where $\mathcal{\hat{Y}}^{i}$ does not entail $\hat{y}^{i,3}\prec\hat{y}^{i,1}$. This is formally defined as:
$P_{\text{TI}}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[R_{MCQA}(V^{i},\{\hat{y}^{i,1},\hat{y}^{i,3}\})=\hat{y}^{i,3}\right].$ (7)
The results, presented in Table 3, demonstrate strong consistency between individual paired caption orders and final caption orders in VLLM predictions.
Figure 7: Overlapping ratios of model predictions under single-frame and full-video inputs for correct, incorrect and overall predictions. Complete Reliance indicates that the VLLM always produces the same response for both video and single-frame inputs.
Image Prior Reliance
Prior research reveals that VLLMs often leverage image priors for reasoning [22, 3], neglecting important spatiotemporal features. This phenomenon is exemplified by a few frames having a dominant influence on response generation. To investigate this bias influences hallucination generation in video-based VLLMs, we first utilized a video summarization algorithm [47] to extract the most salient frame $v^{i}$ from $V^{i}$. Next, we generated VLLM responses on VidHal using $v^{i}$ instead of $V^{i}$ as the the visual input. The influence of image priors is evaluated by identifying overlapping instances where VLLM responses based on the original video $V^{i}$ and the salient frame $v^{i}$ remain consistent, across both correctly and incorrectly predicted orderings. As illustrated in Figure 7, our results highlight that VLLMs exhibit strong reliance on image priors in their predictions. This phenomenon is particularly severe for smaller models such as VideoLLaMA2 (7B).
Figure 10: Qualitative examples of VLLM responses on the caption ordering tasks, for the Attribute, Order and Action aspects.
Figure 8: (Top) Invalid response rates across all models. VLLMs with no invalid responses are grouped under Others. (Bottom) Regurgitation rates of VLLMs on VidHal. Random and Dataset Statistic indicate the regurgitation rates of the random baseline and ground truth answers, respectively. For both metrics, a lower value indicates better model performance.
Figure 9: Hallucination misalignment (HM) scores on VidHal, with Random representing HM scores from the random baseline.
5.4 Further Analysis
Response Stability
The unique structure of the naive caption ordering task, distinct from existing visual understanding tasks, may lead to inadequate model responses due to potential gaps in understanding. To identify these deficiencies, we use the two metrics outlined below:
  • Regurgitation Rate (RR) captures the model’s propensity to consistently generate the same response regardless of input. We define this as the maximum proportion of instances in VidHal where a specific caption order is predicted by the VLLM, across all possible orderings.
  • Invalid Response Rate (IRR) measures the proportion of VLLM responses that fail to provide a valid caption order for the naive caption ordering task.
Figures 8 presents the IRR and RR scores for all evaluated models. Despite being able to formulate responses with the correct structure, many VLLMs tend to produce the same caption order regardless of $V^{i}$, as reflected by their high RR scores. This is particularly evident in models with limited capacity, such as those in the 7B parameter range.
Hallucination Differentiation Sensitivity
We investigate the tendency of VLLMs to favor captions with higher hallucination over those with lower degree in the relative caption ranking task. For two captions with different hallucination levels $j,k$ where $j>k$, we introduce the following metric to quantify such hallucination misalignment cases:
$HM_{j\rightarrow k}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[\mathcal{Y}^{i,j}_{*}\prec\mathcal{Y}^{i,k}_{*}\right].$ (8)
which reflects the proportion of cases in which the VLLM selects the caption with a higher level of hallucination $j$ over $k$. Specifically, we examine three key cases: when the most hallucinatory caption is chosen over both the lower-hallucination and anchor captions, and when the lower-hallucination caption is selected over the anchor caption. These cases are represented by $HM_{3\rightarrow 1}$, $HM_{3\rightarrow 2}$, and $HM_{2\rightarrow 1}$, respectively, with results presented in Figure 9.
Our findings show that advanced VLLMs, such as VideoLLaMA2 (72B), can generally distinguish positive captions from severely hallucinated ones, as reflected by their low $HM_{3\rightarrow 1}$ scores in Figure 9. However, two key observations emerge from our experiments: First, most VLLMs struggle to differentiate the lower hallucinatory caption from the anchor, as evidenced by the gap between $HM_{3\rightarrow 1}$ and $HM_{2\rightarrow 1}$. Second, all models exhibit high $HM_{3\rightarrow 2}$ scores, indicating difficulty in distinguishing between two hallucinatory captions with varying degrees. These results suggests that gaps in nuanced video reasoning may contribute to hallucinatory behavior in VLLMs, a challenge not addressed by existing ⟨positive, hallucinatory⟩-based evaluation methods. [29, 51, 52, 16].
5.5 Qualitative Results
We conducted a qualitative analysis of responses generated by various VLLMs for the caption ordering task, with examples shown in Figure 10. We observe that: I) Relative caption ordering generally guides VLLMs to produce more accurate responses, as evidenced by improvements from naive to relative caption order predictions in most cases. II) Advanced VLLMs exhibit more stable performance across both ordering tasks, reflected by by minimal variation in predictions between the two sub-tasks.
6 Conclusion
In this work, we introduce the VidHal benchmark to address gaps in the video-based hallucination evaluation of VLLMs. VidHal features video instances spanning five temporal aspects. Additionally, we propose a novel caption ordering evaluation task to probe the fine-grained video understanding capabilities of VLLMs. We conduct extensive experiments on VidHal through the evaluation of thirteen VLLMs, exposing their limitations in unexpected hallucination generation. Our empirical results shed light on several promising directions for future work: e.g., incorporating a broader range of temporal features during pretraining and mitigating single-frame priors to enhance temporal reasoning. These advancements will help to address the hallucination problem in video-based VLLMs, enhancing their robustness for real-world video understanding applications.